# Computer Vision-based Adaptive Control for Back Exoskeleton Performance Optimization

### v1

@@ IV-B MobileNetv2 - DINOv2 comparison
As discussed in Section II-B2, we compared two deep learning models, namely MobileNetv2 and DINOv2, primarily used to map input images into meaningful feature spaces. Figure 5-I illustrates the training and validation performance of the two models. Categorizing box weights based on their contents proved challenging, introducing noise into both training processes, particularly due to dataset variability. Nonetheless, the DINOv2-based model demonstrated more stable, reliable, and less overfitting training compared to MobileNetv2-based one, achieving better performance while trained for a lower number of epochs in terms of higher accuracy improving from $90.32\%$ to $94.72\%$ and average precision from $90.23\%$ to $94.83\%$. We found that these improvements were even more pronounced on other datasets, which we provide in the supplementary materials.
@@ IV-C Pipeline Validation
As outlined in Section III-B, during experiments we measure muscular activity to be compared with no-exoskeleton condition, we track output classifications of the pipeline, and ensure that each classification is correct only if the correct output of the classification is provided before the lifting cycle starts. We also collect human feedback on discomfort and preference similarly to what was done for baseline experiments to build validation functions. As shown in Figure 3 (II-a), regarding back muscle activation, we observed unexpected changes under the low weight condition, even though the light assistance selected by the pipeline should be the same as the one provided during the static assistance test. Improvements in the case of adaptive control are registered ($4.90\rightarrow 11.07\%$ for mean and $8.35\rightarrow 19.22\%$ for peak activation), which we attribute to reduced overall effort during the test, and some light weight misclassification leading to stronger support. A strong improvement is registered for the medium and high weight conditions, where adapting the control to the handled weight strongly improves muscular activation reduction compared to static control condition ($9.13\rightarrow 18.89\%$ for mean and $9.62\rightarrow 23.05\%$ for peak activation for medium weight and $12.15\rightarrow 20.68\%$ for mean and $9.99\rightarrow 19.05\%$ for peak activation for heavy weight). We obtain similar results in terms of gradients when combining all of the muscles contributions together. We achieve instead a particularly interesting result when considering leg muscles, which also experience a decrease in peak ($12.35\rightarrow 1.08\%$ for medium weight) and mean ($19.33\rightarrow-2.65\%$ for medium weight) muscle activation when using an adaptive control compared to a static one during lifting. Overall, adaptive control shows strong improvements in EMG reduction across all the muscles. In alignment with baseline tests, D-VRF suggests an increase in perceived discomfort going towards adaptive control, especially marked for medium and heavy weights, while remaining almost unchanged when handling a light weight. This is a result aligned with baseline results, where stronger support led to increased perceived discomfort. P-VRF also aligns with results obtained from the baseline experiments: despite the discomfort slightly increasing, subjects still prefer higher assistance for higher weights, leading to preferring adaptive control compared to static one. Finally, all the T-VRF display gradients pointing towards adaptive control per each of the payload conditions. In conclusion, 83.33% of the subjects agreed in feeling safe letting the exoskeleton autonomously decide which control to select. Figure 6 shows the final validation performances of the pipeline in terms of confusion matrix and accuracies per each subject test and average final value. During real-time tests, the pipeline achieves 85.30%, 76.30% and 86.10% precisions for the light, medium, and heavy weight classes respectively, and an overall 82.41% accuracy. Overall, the pipeline shows balance between overestimations and underestimations of the weights, suggesting that no particular bias is present.
@@ V-A Exoskeleton Optimization Space and Contributions
Results in Section IV-A and Figure 3 (I) indicate that different support modulation strategies can optimize distinct performance metrics. For instance, maintaining high support effectively reduces both peak and mean back muscle activation. However, this approach increases peak activation in leg muscles, particularly during light-weight lifting. While not extensively documented, some research has indicated a comparable pattern, wherein the energy imparted by a non-grounded back exoskeleton seems to distribute throughout the body, possibly reducing back muscle engagement but also, under specific conditions, slightly altering leg loading [41, 42]. Given the negative impact on peak leg muscle activation when using strong support for light weight, an optimal approach for minimizing overall muscular effort would involve high support for reducing mean activation and lower support to mitigate peak activation. Alternatively, considering back and leg muscle activity optimization separately, the best strategy would involve increasing support progressively with weight, from low to strong, ensuring a stable reduction in leg muscle activation. Regarding discomfort, results show that increasing support, regardless of payload, leads to higher perceived discomfort. This finding is consistent with our experience and participants’ feedback, which suggest that greater support exacerbates joint misalignment, increases human-exoskeleton contact forces, and restricts natural movement, all contributing to discomfort. Comments gathered during experiments highlight that strong assistance made bending down more challenging, worsened joint misalignment, and caused localized pressure, particularly at anchoring points such as the thighs and shoulders. Nonetheless, user preference indicates a tendency to favor strong support for heavy loads and low support for lighter ones. This likely stems from a perceived benefit: for short-duration tasks like single-cycle lifting, the exoskeleton’s back muscle relief outweighs the discomfort, whereas for light loads, users do not find the additional support necessary, so they prefer to prioritize discomfort minimization. Once again, the preference function suggests that adaptive support is the most effective strategy. In the case of T-ORF, the support modulation trend is evident across all EMG-ORF contributions, highlighting that maintaining a static support level is suboptimal for maximizing EMG reduction, minimizing discomfort, and aligning with user preferences. Instead, an adaptive control strategy adjusting support from light to strong based on weight would yield the best overall performance. To the best of the authors’ knowledge, this is the only study that considers multiple metrics in the effort to optimize exoskeleton performance.
@@ V-B Payload Estimation Models and Importance of Computer Vision-Based Techniques to Beat Latency
Our results in Section IV-B confirm that payload estimation from transparent boxes remains a challenging task. Despite achieving high accuracy (90.3294.7290.32-94.72%), factors such as varying viewpoints, light reflections, and occlusions caused by box shapes can still lead to misclassification. The MobileNetv2-based model exhibits unstable learning curves, raising concerns about its real-time reliability, even though the final epoch achieves high validation accuracy. Conversely, state-of-the-art Vision Transformer (ViT) architectures like DINOv2 demonstrate more stable training (Figure 5-I) and superior overall performance. While it is less pronounced in DINOv2, both models exhibit some degree of overfitting. However, since our objective is not to develop a zero-shot payload estimation model and industrial environments are typically standardized, training on a dataset collected within the same setting minimizes the risk of encountering vastly different objects, making this overfitting effect acceptable. Feature learning from raw images has been central to deep learning-based computer vision, particularly in CNN applications [43, 44]. However, ViT architectures, especially those trained in an unsupervised manner, have shown superior general feature extraction capabilities, outperforming CNNs in various applications. Our findings further support this, as feature map analysis confirms that DINOv2 extracts higher-quality features, enabling more advanced image detection, segmentation, and depth perception capabilities that standard CNNs, even optimized ones, struggle to achieve. Figure 5-II illustrates how these advantages allow the DINOv2-based model to distinguish foreground from background, filter out irrelevant elements such as floor markings and labels, and demonstrate depth perception crucial for volume and shape estimation. These advanced capabilities result in correct classifications for DINOv2, while MobileNetv2 more frequently misclassifies objects. Real-time evaluations confirm strong performance. As shown in Figure 6, the system achieves a mean accuracy of 82.4182.41%, with subject-specific accuracies closely distributed around this mean, indicating that our approach effectively mitigates the subject-specific variability found in other payload estimation methods [23]. Accuracy drops for subjects 3 and 8 are primarily due to communication conflicts between the cloud computing platform and the embedded Raspberry Pi, arising from interference between the embedded computer and the EMG wireless sensor system, both operating at the same frequency. While this highlights a limitation in our pipeline, it also suggests that improved system design could further enhance stability and accuracy. Additionally, the confusion matrix shows that misclassifications are evenly distributed around the diagonal, indicating a well-balanced model without inherent bias. All correctly classified transitions occurred before the lifting cycle began, ensuring prompt detection and support modulation. To the best of our knowledge, this is the only technique available to date that enables payload estimation and support modulation in advance, maximizing exoskeleton effectiveness and benefits. Our results in Figure 3 (II), particularly in terms of EMG-VRF, strongly support this claim.
@@ V-C The Importance of Adaptive vs Static Control
Results in Section IV-C indicate that, except when optimizing solely for discomfort, adaptive assistance is consistently preferred. Figure 3 (II) shows that adaptive control enhances EMG reduction across all contributions, even in cases where support levels should theoretically be the same, such as for light weights, where both static and adaptive assistance provide low support. An interesting finding is the increase in peak and mean leg muscle activation observed during experiments, particularly for the light-weight static assistance condition, which slightly deviates from baseline results. In addition to the pipeline’s non-perfect classifications, we attribute this to differences in test protocols: while baseline tests involved controlled, standardized stoop lifting, validation tests allowed participants to move freely, lifting at their preferred speed and with their preferred style. As a result, some subjects performed a mixed stoop-squat lifting motion, leading to increased leg muscle activation. Video analysis and EMG data confirm this effect, particularly for subjects 2 and 8. This effect suggests that adaptive control allows subjects to lift weights more comfortably without relying on leg muscles, leading to greater muscle activity reduction in both the back and legs. For payload increasing static support does not substantially enhance muscle activity reduction, whereas adaptive support does, as expected, by dynamically adjusting support based on payload. This trend is evident across all contributions. Regarding discomfort, the VRF results align with expectations, showing an increase under adaptive control. This supports the baseline findings that higher support levels lead to greater discomfort, as adaptive control adjusts assistance based on payload, while static support maintains low assistance regardless of weight. Similarly, the preference VRF indicates that most subjects favored adaptive control, as it provided greater support for heavier loads while minimizing discomfort for lighter ones by adapting to low assistance. When considering all factors together, the T-VRF results reveal a clear trend favoring adaptive control for performance optimization, validating the hypotheses drawn from baseline experiments. Eventually, our findings support the claim that adaptive control is necessary for performance optimization, suggesting a preference for active exoskeletons over passive ones.
@@ V-D Limitations & Future Work
Our pipeline shows promising results in real-time detection and validation, demonstrating that adaptive control can significantly improve performance. However, it has limitations that may hinder adoption in industrial settings. A key constraint is its reliance on transparent boxes, which aid detection and contextual inference. Estimating payloads in opaque boxes is more challenging, as even humans struggle without prior knowledge. In environments with known object weights (e.g., cataloged), the payload estimation layer can be bypassed, but alternative approaches, such as sensor fusion, are needed when a catalog is unavailable. Future work will explore continuous payload estimation by integrating EMG/FSM data for refinement in critical scenarios. Another limitation is that our method has only been tested in a controlled laboratory setting with young subjects. Future research should evaluate the pipeline in diverse industrial environments with real workers. Additionally, our pipeline currently runs on the cloud, which aligns with the Industry 5.0 paradigm but limits system independence and introduces potential real-time issues, such as communication delays. Future work should focus on optimizing hardware and software to enable embedded execution without increasing the exoskeleton’s bulk.
@@ VI Conclusion
In this research, we addressed two key questions: identifying optimization spaces for back exoskeletons to determine optimal operating areas and exploring control adaptation strategies to enhance performance through exoskeleton context awareness. We established a two-dimensional optimization space using Optimization and Validation Representation Functions based on EMG reduction, perceived discomfort, and user preference. Experiments with 12 subjects revealed optimal operating regions, demonstrating that support modulation is essential for performance optimization. Building on these insights, we developed a computer vision-based adaptive control pipeline. Unlike previous approaches, our method is subject-independent, does not increase exoskeleton bulk, and minimizes latency by estimating payloads before lifting. Integrated into the exoskeleton control and validated with 12 subjects, the pipeline demonstrated strong real-time accuracy and effectiveness across performance metrics. Our findings highlight the importance of adaptive control in optimizing exoskeleton performance and suggest that active exoskeletons could be preferred over passive/static ones to achieve the necessary adaptability. These results pave the way for further research in back exoskeleton optimization, context awareness, and high-level control adaptation.
@@
Conflict of Interests
The authors declare that M.G. and F.B. hold shares in AGADE srl and AllyArm srl, and there were no further competing interests in the development of this research.

### v2

@@ IV-B MobileNetv2 - DINOv2 comparison
As discussed in Section II-B2, we compared two deep learning models, namely MobileNetv2 and DINOv2, primarily used to map input images into meaningful feature spaces. Figure 6-I illustrates the offline training and validation performance of the two models. Categorizing box weights based on their contents proved challenging, introducing noise into both training processes, particularly due to dataset variability. Nonetheless, the DINOv2-based model demonstrated more stable, reliable, and less overfitting training compared to MobileNetv2-based one, achieving better performance while trained for a lower number of epochs in terms of higher accuracy improving from $90.32\%$ to $94.72\%$ and average precision from $90.23\%$ to $94.83\%$. We found that these improvements were even more pronounced on other datasets, which we provide in the Supplementary Materials.
@@ IV-C Pipeline Validation
As outlined in Section III-B, during experiments we measure muscular activity to be compared with no-exoskeleton condition, we track output classifications of the pipeline, and ensure that each classification is correct only if the correct output of the classification is provided before the lifting cycle starts. We also collect human feedback on discomfort and preference similarly to what was done for baseline experiments to build validation functions. As shown in Figure 4 (II-a), regarding back muscle activation, we observed unexpected changes under the light weight condition, even though the low assistance selected by the pipeline should be the same as the one provided during the static assistance test. We registered improvements in the case of adaptive control ($4.90\rightarrow 11.07\%$ for mean and $8.35\rightarrow 19.22\%$ for peak activation), which we attribute to reduced overall effort during the test, and some light weight misclassification leading to stronger support. We registered a strong improvement for the medium and heavy weight conditions, where adapting the control to the handled weight strongly improves muscular activation reduction compared to static control condition ($9.13\rightarrow 18.89\%$ for mean and $9.62\rightarrow 23.05\%$ for peak activation for medium weight and $12.15\rightarrow 20.68\%$ for mean and $9.99\rightarrow 19.05\%$ for peak activation for heavy weight). We obtain similar results in terms of gradients when combining all of the muscles contributions together. We achieve instead a particularly interesting result when considering leg muscles, which also experience a decrease in peak ($12.35\rightarrow 1.08\%$ for medium weight) and mean ($19.33\rightarrow-2.65\%$ for medium weight) muscle activation when using an adaptive control compared to a static one during lifting. Overall, adaptive control shows strong improvements in EMG reduction across all the muscles. In alignment with baseline tests, D-VRF suggests an increase in perceived discomfort going towards adaptive control, especially marked for medium and heavy weights, while remaining almost unchanged when handling a light weight. This is a result aligned with baseline results, where stronger support led to increased perceived discomfort. P-VRF also aligns with results obtained from the baseline experiments: despite the discomfort slightly increasing, subjects still prefer higher assistance for heavier weights, leading to preferring adaptive control compared to static one. Finally, all the T-VRF display gradients pointing towards adaptive control for each of the payload conditions. In conclusion, 83.33% of the subjects agreed in feeling safe letting the exoskeleton autonomously decide which control to select.
Concerning the CVAC performances, using video recordings of the experiments, we evaluated the accuracy of the object detection and selection layer described in Section II-B. The method performed robustly, achieving 98.13% accuracy in real-time trials. While Figure 7 shows the final validation performances of the whole CVAC pipeline in terms of confusion matrix and accuracies per each subject test and average final value. During real-time tests, the pipeline achieves 85.30%, 76.30% and 86.10% precisions for the light, medium, and heavy weight classes respectively, and an overall 82.41% accuracy. Overall, the pipeline shows balance between overestimations and underestimations of the weights, suggesting that no particular bias is present.
@@ V-A Exoskeleton Optimization Space and Contributions
Results in Section IV-A and Figure 4 (I) indicate that different support modulation strategies can optimize distinct performance metrics. For instance, maintaining high support effectively reduces both peak and mean back muscle activation. However, this approach increases peak activation in leg muscles, particularly during light-weight lifting. While not extensively documented, some research has indicated a comparable pattern, wherein the energy imparted by a non-grounded back-support exoskeleton seems to distribute throughout the body, possibly reducing back muscle engagement but also, under specific conditions, slightly altering leg loading [13], [21]. Given the negative impact on peak leg muscle activation when using strong support for light weight, an optimal approach for minimizing overall muscular effort would involve high support for reducing mean activation and lower support to mitigate peak activation. Alternatively, considering back and leg muscle activity optimization separately, the best strategy would involve increasing support progressively with weight, from low to strong, ensuring a stable reduction in leg muscle activation. Regarding discomfort, results show that increasing support, regardless of payload, leads to higher perceived discomfort. This finding is consistent with our experience and participants’ feedback, which suggest that greater support exacerbates joint misalignment, increases human-exoskeleton contact forces, and restricts natural movement, all contributing to discomfort. Comments gathered during experiments highlight that strong assistance made bending down more challenging, worsened joint misalignment, and caused localized pressure, particularly at anchoring points such as the thighs and shoulders. Nonetheless, user preference indicates a tendency to favor strong support for heavy loads and low support for lighter ones. This likely stems from a perceived benefit: for short-duration tasks like single-cycle lifting, the exoskeleton’s back muscle relief outweighs the discomfort, whereas for light loads, users do not find the additional support necessary, so they prefer to prioritize discomfort minimization. Once again, the preference function suggests that adaptive support is the most effective strategy. In the case of T-ORF, the support modulation trend is evident across all EMG-ORF contributions, highlighting that maintaining a static support level is suboptimal for maximizing EMG reduction, minimizing discomfort, and aligning with user preferences. Instead, an adaptive control strategy adjusting support from light to strong based on weight would yield the best overall performance. Although it is intuitive that heavier payloads require greater assistance, the quantitative, multi-metric landscape that determines the optimal level of assistance balancing multiple metrics has not been characterized in prior work. In this study, we provide the first systematic and empirical analysis of this landscape and show, through the results in Figure 5, that it follows a non-linear, subject-independent exponential trend that cannot be inferred from intuition alone. To the best of our knowledge, this is the first study to jointly consider multiple metrics in order to derive an optimized assistance profile for back exoskeletons.
@@ V-B Payload Estimation Models and Importance of Computer Vision-Based Techniques to Beat Latency
Our results in Section IV-B confirm that payload estimation from transparent boxes remains a challenging task. Despite achieving high accuracy (90.3294.7290.32-94.72%), factors such as varying viewpoints, light reflections, and occlusions caused by box shapes can still lead to misclassification. The MobileNetv2-based model exhibits unstable learning curves, raising concerns about its real-time reliability, even though the final epoch achieves high validation accuracy. Conversely, state-of-the-art Vision Transformer (ViT) architectures like DINOv2 demonstrate more stable training (Figure 6-I) and superior overall performance. While it is less pronounced in DINOv2, both models exhibit some degree of overfitting. However, since our objective is not to develop a zero-shot payload estimation model and industrial environments are typically standardized, training on a dataset collected within the same setting minimizes the risk of encountering vastly different objects, making this overfitting effect acceptable. Feature learning from raw images has been central to deep learning-based computer vision, particularly in CNN applications [42], [30]. However, ViT architectures, especially those trained in an unsupervised manner [4, 8], have shown superior general feature extraction capabilities, outperforming CNNs in various applications. Our findings further support this, as feature map analysis confirms that DINOv2 extracts higher-quality features, enabling more advanced image detection, segmentation, and depth perception capabilities that standard CNNs, even optimized ones, struggle to achieve. Figure 6-II illustrates how these advantages allow the DINOv2-based model to distinguish foreground from background, filter out irrelevant elements such as floor markings and labels, and demonstrate depth perception crucial for volume and shape estimation. These advanced capabilities result in correct classifications for DINOv2, while MobileNetv2 more frequently misclassifies objects. Real-time evaluations confirm strong performance. While our object detection and selection layer achieved 98.13% accuracy, we acknowledge that it may still benefit from additional hyperparameter tuning in more challenging environments. As shown in Figure 7, the end-to-end system achieves a mean accuracy of 82.4182.41%, with subject-specific accuracies closely distributed around this mean, indicating that our approach effectively mitigates the subject-specific variability found in other payload estimation methods [44]. Accuracy drops for subjects 3 and 8 are primarily due to communication conflicts between the cloud computing platform and the embedded Raspberry Pi, arising from interference between the embedded computer and the EMG wireless sensor system, both operating at the same frequency. While this highlights a limitation in our pipeline, it also suggests that improved system design could further enhance stability and accuracy. Additionally, the confusion matrix shows that misclassifications are evenly distributed around the diagonal, indicating a well-balanced model without inherent bias. All correctly classified transitions occurred before the lifting cycle began, ensuring prompt detection and support modulation. To the best of our knowledge, this is the only technique available to date that enables payload estimation and support modulation in advance, maximizing exoskeleton effectiveness and benefits. Our results in Figure 4 (II), particularly in terms of EMG-VRF, strongly support this claim.
@@ V-C The Importance of Adaptive vs Static Control
Results in Section IV-C indicate that, except when optimizing solely for discomfort, adaptive assistance is consistently preferred. Figure 4 (II) shows that adaptive control enhances EMG reduction across all contributions, even in cases where support levels should theoretically be the same, such as for light weights, where both static and adaptive assistance provide low support. An interesting finding is the increase in peak and mean leg muscle activation observed during experiments, particularly for the light-weight static assistance condition, which slightly deviates from baseline results. In addition to the pipeline’s non-perfect classifications, we attribute this to differences in test protocols: while baseline tests involved controlled, standardized stoop lifting, validation tests allowed participants to move freely, lifting at their preferred speed and with their preferred style. As a result, some subjects performed a mixed stoop-squat lifting motion, leading to increased leg muscle activation. Video analysis and EMG data confirm this effect, particularly for subjects 2 and 8. This effect suggests that adaptive control allows subjects to lift weights more comfortably without relying on leg muscles, leading to greater muscle activity reduction in both the back and legs. For payload increasing static support does not substantially enhance muscle activity reduction, whereas adaptive support does, as expected, by dynamically adjusting support based on payload. This trend is evident across all contributions. Regarding discomfort, the VRF results align with expectations, showing an increase under adaptive control. This supports the baseline findings that higher support levels lead to greater discomfort, as adaptive control adjusts assistance based on payload, while static support maintains low assistance regardless of weight. Similarly, the preference VRF indicates that most subjects favored adaptive control, as it provided greater support for heavier loads while minimizing discomfort for lighter ones by adapting to low assistance. When considering all factors together, the T-VRF results reveal a clear trend favoring adaptive control for performance optimization. As noted in the experiments section, we intentionally adopted distinct protocols to test whether optimal performance trends observed in standardized conditions persist in more dynamic environments. The results confirm this, as adaptive control remained preferable to the static strategy even under more variable lifting conditions. Eventually, our findings support the claim that adaptive control is necessary for performance optimization, suggesting a preference for active exoskeletons over passive ones.
@@ V-D Limitations & Future Work
Our pipeline shows promising results in real-time detection and validation, demonstrating that adaptive control can significantly improve performance. However, it has limitations that may hinder adoption in industrial settings. A key limitation of this study is that payload estimation was tested primarily with transparent boxes, which simplify detection and inference. Estimating weights for opaque or irregularly shaped objects, common in industrial settings, remains challenging. While opaque boxes with attached payload information (e.g., QR codes or labels) can be handled by fine-tuning DINOv2, estimating weights for irregular or opaque objects without explicit cues remains challenging and may benefit from additional sensing, such as sensor fusion or EMG/FSM data. Human motor control achieves robust performance through a combination of predictive (visual) and reactive (tactile) mechanisms. In this work, we isolate and validate the predictive visual component, while future extensions could integrate it with real-time haptic feedback to achieve high-performance, dual-modality exoskeleton assistance, mirroring the reliability of the human sensorimotor system. Furthermore, incorporating dual heterogeneous sensors and sensor-fusion strategies would provide hardware-level redundancy and fault-tolerant monitoring, which are necessary to meet safety requirements for real-world deployment. Additionally, our current pipeline runs on the cloud, which is suitable for the experimental validation of our framework and aligns with the Industry 5.0 paradigm. However, it also introduces certain limitations, such as reduced system autonomy and potential real-time issues caused by communication delays. Future work will therefore focus on software optimization and the integration of edge-computing capabilities to ensure reliable, low-latency performance even in environments with unstable network connectivity. Ultimately, our goal is to enable fully embedded execution without increasing the bulk or weight of the exoskeleton. Another limitation is that our method was tested only in a controlled lab with young subjects and a three-level assistance scheme. While suitable for the experimental validation, future research efforts should be put toward continuous assistance modulation, improved payload estimation, and testing with more diverse objects. Advancing beyond the proof of concept will require evaluations in real industrial environments with varied user profiles to assess scalability, robustness, and latency performance.
@@ VI Conclusion
In this research, we addressed two key questions: identifying optimization spaces for back-support exoskeletons to determine optimal operating areas and exploring control adaptation strategies to enhance performance through exoskeleton context awareness. We established a two-dimensional optimization space using Optimization and Validation Representation Functions based on EMG reduction, perceived discomfort, and user preference. Experiments with 12 subjects revealed optimal operating regions, demonstrating that support modulation is essential for performance optimization. Building on these insights, we developed a computer vision-based adaptive control pipeline. Unlike previous approaches, our method is subject-independent, does not increase exoskeleton bulk, and minimizes latency by estimating payloads before lifting. Integrated into the exoskeleton control and validated with 12 subjects, the pipeline demonstrated strong real-time accuracy and effectiveness across performance metrics. Our findings highlight the importance of adaptive control in optimizing exoskeleton performance and suggest that active exoskeletons could be preferred over passive/static ones to achieve the necessary adaptability. These results pave the way for further research in back exoskeleton optimization, context awareness, and high-level control adaptation.